Compute the magnitude of the faintest object observed with a 1 m optical telescope.
1
Expert's answer
2019-05-30T09:05:31-0400

Solution. More light enters the telescope, the fainter objects can be viewed. Limiting magnitude of a telescope can be found according the following formula: 


"M_L=2.5^m+5\times lg D"

where D is diameter of the main lens or mirror of the telescope in millimeters. Therefore


"M_L=2.5^m+5\times lg 1000=17.5^m"

Answer.

"17.5^m"